### v1

@@
1 Introduction
Feature Interaction, one of the key challenges in predictive modeling, is the behavior of datasets whose two or more independent variables have a common influence on their dependent variable. Mathematically, it refers to operations between features in a conjunctive manner, such as multiplication, rather than treating features in isolation or through simple linear combinations. For example, if we have a dataset generated by the equation $y=x_{1}x_{2}+x_{3}$, then we say that $x_{1}$ and $x_{2}$ interact with each other in the model of $y$.
Modeling interactions can significantly improve the accuracy of predictions and reduce the complexity of models by augmenting datasets with interaction features for the models to learn from. In the best-case scenario, even linear models can be used effectively. For example, it is easier to approximate the previous function by including interaction terms $x_{1}x_{2}$ and $x_{3}$ in the dataset, rather than using more complex models with only the features $x_{1},x_{2}$ and $x_{3}$ with more complex models. Although this approach is not straightforward, it improves the interpretability of downstream tasks.
Applications of feature interaction modeling span a wide range of domains, particularly in recommendation systems, where it plays a critical role in improving predictions. The click-through rate (CTR) problem in recommendation systems is an example of an application highly based on feature interactions. For example, the age of the user and the type of product may interact in different ways on various platforms. Younger users may prefer to buy fashion products on one platform, while older users might lean towards formal products on all platforms.
Although many models, including traditional machine learning methods like decision trees, are capable of modeling feature interactions through their hierarchical structures. For example, we may consider that, in the context of meal preferences, a decision tree could represent the interaction between a type of wine and a kind of food (steak or fish) by placing the wine-related node as a descendant of the food-related node. Although this structure provides a clear representation of feature interactions, these models face significant limitations when dealing with complex or high-dimensional interactions. As the complexity of the interactions increases, the trees often become too large, leading to reduced interpretability.
Hence, we focus on graph-based methods which are more expressive in interpretability with edges between nodes. Specifically, graph-based methods represent data using graphs and model interactions through Graph Neural Networks (GNNs). Many graph-based methods have been developed to model feature interactions. Feature graphs, where nodes represent features, play an important role in GNN-based models for modeling interactions. However, a major challenge arises because most methods rely on a complete graph, where every possible interaction, i.e., every pair of features, is included. In addition, no insight and analysis about structures of feature graphs concerning features interactions are studied to make us understand the effect of graph structure to model it. One of the key questions that arises is the following: What characteristics should a feature graph have to optimally model interactions, instead of relying only on complete graphs?
Understanding the relationship between feature graph structures and their ability to model feature interactions is essential to building efficient and interpretable representations. Using a complete graph often introduces unnecessary complexity and increases computational overhead. Identifying and selectively including only the most meaningful edges in a feature graph can be considered one of the feature selection techniques for graph models that can avoid the overfitting issue.
In this work, we attempt to discover this insight, focusing on pairwise interactions, interactions between two features. We do not propose a new method; instead, we utilize existing methods in our experiments as tools for investigation. Our key contributions are as follows.
@@ 1 Introduction
    Theoretical Foundation for Sparse Feature Graphs: We establish a theoretical foundation for sparse feature graphs using the Minimum Description Length (MDL) principle. By framing the feature graph selection problem in the context of MDL, we prove that retaining only necessary interaction edges minimizes the description length, making it a better choice than using complete graphs. (Section 5)
@@
    Insights into Feature Graph Structures: We provide insights into the structures of feature graphs pertaining the given feature interactions instead of using complete graphs to be used as input graphs for GNN models. Our experiment result shows that the edges corresponding to the pair of interacting features are significant and must remain in the feature graphs to maintain the prediction performance. (Section 7)
@@
    Empirical Validation on Real-World Datasets: We empirically show that, on real-world datasets, the feature graphs constructed by the pairwise interaction can improve the prediction performance compared with baselines including tree-based algorithms and linear-feature-based algorithms. (Section 8)
2 Preliminary
@@
Table 1: All related terminology
@@
For graphs, we denote a tuple $G=(V,E,A,X)$ of nodes $V$, a set of edges $E$, an adjacency matrix $A$ and an embedding matrix $X$. For any node $v\in V$, we denote by $\mathcal{N}(v)$ the set of nodes adjacent to $v$, called the neighborhood of the node $v$. If only the structure of a graph is considered, we may use only $G(V,E)$.
Since we represent the interaction between features by graph edges, we model instances with the graphs whose nodes correspond to features of a dataset. According to Fi-GNN [1], this graph is called a feature graph. (see the example in Figure 2) Mathematically, given a dataset of features $x_{1},\dots,x_{d}$, a feature graph structure is a graph $G(V,E)$ where $V=\{v_{1},\dots,v_{d}\}$.
There are many tools that can be used to process graphs. Among these GNN models have gained significant attention due to their effectiveness in handling complex, non-Euclidean data. GNNs are a category of neural models specifically designed to operate on graph data structures, where the order of nodes should not influence the computation of predicted values, a property referred to as “permutation invariance.” There are various approaches to implement GNNs, but one of the most prominent paradigms is based on “message passing.”
@@
where $\bigoplus$ represent a differentiable, permutation-invariant function, such as summation, averaging, or maximum operation, while $\gamma$ and $\phi$ denote differentiable functions, often implemented as neural network layers. In Equation 2, the model updates the node information using both existing information (from the previous layer) and information shared among neighboring nodes (messages), which is computed as Equation 1. The latter equation pertains to the aggregation of messages from neighboring nodes.
@@
Our problem can be framed as the selection of models, specifically the representation of graphs. MDL [3, 4] is a paradigm of model selection that attempts to find the model that minimizes the number of bits that represent information. This information may include the complexity of models and the performance score of the model on given data.
In our work, we use MDL to study the model whose complexity is minimized while maintaining the well-performed accuracy of the prediction. It is the two-stage code description approach [5]. Specifically, it encodes the data $D$ with the description length $\mathcal{L}_{\mathcal{H}}(D)$ by first encoding a hypothesis $H$ from the set of given hypotheses $\mathcal{H}$ followed by encoding $D$ with the help of $H$. The objective is to minimize the total description length as follows:
@@
3 Related Work
3.1 Feature Interaction Modeling
@@
From the perspective of machine learning, many models were proposed to model feature interactions, especially variants of factorization models inspired by FM [7]. FM relies on the inner product between the embedding of the feature-value motivated by the factorization of matrices. Several variants of FM have been proposed to fill the gaps. FFM [8] takes into account the embedding of feature field (i.e. columns in a table of data). AFM [9] considers the weight of the interaction by adding attention coefficients to its model.
@@
The attention mechanisms [12] are the method designed to model the importance between different components. In AFM, the attention mechanism was applied because the authors believe that all interactions should not be treated equally likely with respect to prediction values. Attention allows the model to weight feature interactions differently based on their importance. After introducing the attention mechanism to interaction tasks, many works use attention to model the interaction of features based on the factorization machine framework [13, 14, 15, 16, 17].
One of the crucial challenges of interactions in traditional machine learning models is the explainability of which features have interaction, although some of them outperform in prediction performance. So we need an explicit representation that can correspond to pairwise interaction, where a graph is a representation that can fill this gap. In addition, GNN is a tool that can be used to learn information from graphs.
3.2 Graph Neural Networks for Feature Interactions
GNN are becoming more interesting for use in feature interaction problems. Since feature interactions can be seen as relationships between features, most of the works rely on feature graphs for each individual instance to represent the interactions via aggregation of representation from neighbor nodes.
Motivated by the click-through rate (CTR) prediction, which is believed to influence the rate of clicking on a product, such as age and gender, many works attempted to model this prediction by using GNN applied feature graphs. To the best of our knowledge, Fi-GNN [1] is the first work to use GNN to model the interaction of features in the feature graph for the prediction of CTR. It applied the field-aware embedding layer to compute the latent representation of nodes of features before feeding the feature graph with the computed representations into a stack of message passing layers.
@@ 3.2 Graph Neural Networks for Feature Interactions
All mentioned works chronically improve the methods of feature interaction learning. However, no work provides information about the effect of the construction of feature graphs on the learning capacity. Most of the works utilize attention mechanisms or edge weights to let models learn graph structures via these parameters.
4 Problem Formulation: Pairwise Interaction Feature Graph Problem
An important open research question in graph representation for data without explicit connections is that What is a suitable graph structure? While current research highlighted the potential of feature graphs, a feature graph that is too dense, for example, a complete graph, may lead to learning issues, as shown and discussed in previous work [1, 22, 23] and in Section 7.5. Thus, finding an optimal feature graph structure for any particular prediction task is the common goal of the community. Consequently, we can formulate the “feature-graph-finding” problem as follows:
@@ 4 Problem Formulation: Pairwise Interaction Feature Graph Problem
where $\text{GNN}(\ \cdot\ ;G)$ is a given GNN model trained for the feature graph $G$, and $\ell$ is an arbitrary but fixed loss function for the prediction task. The above task is not traceable as there are exponentially many possible graph structures ($\mathcal{O}(2^{n^{2}})$.) Many works attempted to solve this problem using soft learnable interactions [1, 19] or hard interactions through reinforcement learning [21]. However, they did not provide any insight into the characteristics of a suitable feature graph structure related to the interaction between features.
@@ 5.1 Correspondence Problem
Multiple graphs can infer the same expression considering the perspective of connected components. For example, the expression $x_{0}x_{1}x_{2}+x_{3}x_{4}+x_{5}+x_{6}$ corresponds to the graphs $E_{1}=\{\{0,1\},\{0,2\},\{1,2\},\{3,4\}\}$ and also $E_{2}=\{\{0,1\},\{1,2\},\{3,4\}\}$. Consequently, we use equivalence classes of graphs (based on connectedness) instead of individual graphs. To this end, we define an equivalence class of feature graphs.
@@
5.2 MDL for graph representation of pairwise interaction
@@
Problem 1 MDL Pairwise Interaction Graph Problem******
@@
In this problem, we denote $\mathcal{L}(S,f_{G})=\mathcal{L}(S|f_{G})+\mathcal{L}(f_{G})$, where
@@
The theoretical results shown in Section 5.1 are just to ensure us that we can confide in learning feature graph structures to infer an expression containing feature interactions. In other words, given any function containing feature interactions of at most one occurrence for each feature, there always exist feature graph structures that can infer backward to the function and vice versa. Therefore, to find such graph structures, we need some insight into graph construction related to feature interactions.
Consequently, we reformulate these observations into more general arguments. Generally speaking from empirical results in Section 7.2, it is better to maintain all the edges of the interaction in the feature graph to maximize the learning interaction in a given dataset. Also, it is better to prune non-interaction edges out as much as possible to avoid fitting noise information. Finally, these results are supported by Proposition 5.
In addition, we also show the empirical result about using complete feature graphs to learn from larger datasets in Subsection 7.5. Although the result in Subsection 7.2 cannot infer the limitation of complete graph obviously, the result in Subsection 7.5 reveals that the complete feature graphs are worse when datasets are larger in feature size. Finally, we can show that, in the manner of MDL which is a framework of model selection, the complete feature graph is worse than any feature graphs containing all interaction edges. This tells us that it is not a good idea to use the complete feature graph to learn from datasets. Not only learning hidden noise, but it also face with the computation and memory issues. In the next section, we will focus on the empirical results in more details.
6 Methodology
@@
Our experiment provides empirical evidence regarding the relationship between feature graph structures and pairwise interactions in datasets. We train and evaluate several feature graph structures for each dataset and report the results. However, due to the combinatorial number of possible feature graph structures and the stochastic nature of the learning algorithm, we adopt a strategy of generating random edge sets to explore various graph configurations, as illustrated in Figure 2.
6.1 Datasets
Real-world datasets rarely provide information about interactions between features, which disables our analysis. To this end, we use synthesized datasets to control interaction characteristics in datasets such as the number of interaction pairs of features, strength of interactions, or kinds of interactions. We assume that there is only one mode of interaction per feature graph. If real-world datasets had many interaction modes, we could apply weighted ensemble feature graphs proposed Li et al. [17]
@@ 6.3 GNN Architectures
As Fi-GNN does, this work adopts the idea from Fi-GNN, but simpler architecture. We use TransformerConv [24] which is a message-passing layer based on the attention mechanism whose feature interaction learning capability is shown by many works. For the computation, we start from the initial node embedding $H^{(0)}$ defined in Equation 9. Then a feature graph with the node embedding matrix is fed to TransformerConv with BatchNorm:
@@
where $H^{(l)}$ is an intermediate node embedding matrix at the layer $l$. After applying all message passing layers, we calculate the average pooling of node embeddings from the last GNN layer, and we use this vector as a graph embedding which is fed into the projection layer to get prediction:
@@
where $H^{(L)}$ is a node embedding matrix from the last GNN layer and $H^{(L)}[i]$ is the row $i$ of the matrix which corresponds to the node embedding of the node $i$ and $b$ is a biased term. Equation 11 is the calculation of mean pooling layer and Equation 12 is the final projection to calculate the prediction value.
Equation 10 is a bit different from the original work [24] which uses LayerNorm after aggregation of message like as the original transformer in NLP. In the first phase of experiment, we found that the GNN model with LayerNorm cannot capture interaction behavior while BatchNorm can.
In training the GNN model, we adopt an usual supervised learning approach, where the loss function measures the difference between the predicted value, $\hat{y}$, and the true target value, $y$. Specifically, we use the mean squared error (MSE) loss, which is defined as

### v2

@@
Abstract
Feature interaction is crucial in predictive machine learning models, as it captures the relationships between features that influence model performance. In this work, we focus on pairwise interactions and investigate their importance in constructing feature graphs for Graph Neural Networks (GNNs). We leverage existing GNN models and tools to explore the relationship between feature graph structures and their effectiveness in modeling interactions. Through experiments on synthesized datasets, we uncover that edges between interacting features are important for enabling GNNs to model feature interactions effectively. We also observe that including non-interaction edges can act as noise, degrading model performance. Furthermore, we provide theoretical support for sparse feature graph selection using the Minimum Description Length (MDL) principle. We prove that feature graphs retaining only necessary interaction edges yield a more efficient and interpretable representation than complete graphs, aligning with Occam’s Razor. Our findings offer both theoretical insights and practical guidelines for designing feature graphs that improve the performance and interpretability of GNN models.
Index Terms:
Feature interaction, Graph neural networks
I Introduction
One of the key challenges in predictive modeling is the behavior of data whose two or more independent variables exhibit an interaction effect on their dependent variable – we refer to this as Feature Interaction [1]. The interaction refers to conjunctive operations, such as multiplication, rather than treating features in isolation or through simple linear combinations. For example, if we have a dataset generated by the equation $y=x_{1}x_{2}+x_{3}$, then we say that $x_{1}$ and $x_{2}$ interact with each other in the model of $y$. Applications of feature interaction modeling span a wide range of domains, particularly the click-through rate (CTR) problem, where they play a critical role in improving predictions [2, 3, 4]. For example, the user’s age and product type may interact differently on various platforms. Younger users may prefer to buy fashion products (young and fashionable), while older users might lean towards formal products (old and formal).
@@ I Introduction
Modeling these interactions between variables has been a long-standing research problem in machine learning. Traditionally, these interactions were captured through manual feature engineering, where practitioners explicitly crafted cross-features based on domain knowledge [5]. Powered by the theory for the partial dependence decomposition, Friedman’s H statistic [6] is a method to learn the dependence between features. While effective in some cases, manual approaches are labor-intensive, unscalable, and prone to missing complex or subtle interactions [1]. To address these limitations, models such as the Cross Field Network (CFN) [7] were developed to automatically learn bounded-degree feature interactions, particularly in click-through rate prediction tasks. Recently, graph neural networks have emerged as a powerful framework for modeling interactions through graph-structured representations. In this setting, feature graphs, where nodes represent features, play an important role in GNN-based models for modeling interactions [8]. Figure 1 illustrates examples of feature graphs when a set of features in a dataset is given.
While feature graphs are more expressive in representing feature interactions with edges between nodes, a major challenge arises when we do not know the interactions in advance. Thus, most methods rely on searching or learning from a complete graph, which includes every possible interaction, i.e., every pair of features, to avoid missing any potentially important interaction [9]. For example, Table2Graph [10] formulates the construction of a unified feature graph as a reinforcement learning problem, where the graph’s edge weights are optimized using the reinforce algorithm. In addition, Fi-GNN [9] takes advantage of the robust expressive capabilities of feature graphs to represent the multi-field features in a feature graph structure where different fields can interact through edges. Although these show the potential applications of feature graphs, a key theoretical question remains: what characteristics should a feature graph have to optimally model interactions? Understanding the relationship between feature graph structures and their ability to model feature interactions is essential to building efficient and effective algorithms for manually constructing feature graphs when interactions are known or learning feature graphs when interactions are unknown.
In this work, instead of proposing another feature graph algorithm, we attempt to discover the insights related to the feature graph structures and feature interactions, focusing on pairwise interactions. Our key contributions are as follows.
@@
    Theoretical Foundation for Sparse Feature Graphs: We establish a theoretical foundation for sparse feature graphs using the Minimum Description Length (MDL) principle. By framing the feature graph selection problem in the context of MDL, we prove that retaining only necessary interaction edges minimizes the description length, making it a better choice than using complete graphs. (Section V)
@@
    Insights into Feature Graph Structures: We provide insights into the structures of feature graphs pertaining the given feature interactions instead of using complete graphs to be used as input graphs for GNN models. Our experiment result shows that the edges corresponding to the pair of interacting features are significant and must remain in the feature graphs to maintain the prediction performance. (Section VII)
@@
    Empirical Validation on Real-World Datasets: We empirically show that, on real-world datasets, the feature graphs constructed by the pairwise interaction can improve the prediction performance compared with baselines, including tree-based algorithms and linear-feature-based algorithms. (Section VIII)
II Preliminary
@@
TABLE I: All related terminology
@@
For graphs, we denote a tuple $G=(V,E,A,X)$ of nodes $V$, a set of edges $E$, an adjacency matrix $A$ and an embedding matrix $X$. For any node $v\in V$, we denote by $\mathcal{N}(v)$ the set of nodes adjacent to $v$, called the neighborhood of the node $v$. If only the structure of a graph is considered, we may use only $G(V,E)$. Since we represent the interaction between features by graph edges, we model instances with the graphs whose nodes correspond to features of a dataset. According to Fi-GNN [9], this graph is called a feature graph. (see the example in Figure 2) Mathematically, given a dataset of features $x_{1},\dots,x_{d}$, a feature graph structure is a graph $G(V,E)$ where $V=\{v_{1},\dots,v_{d}\}$.
There are many tools for processing graphs. GNN models have gained significant attention due to their effectiveness in handling complex, non-Euclidean data. GNNs are a category of neural models specifically designed to operate on graph data structures. In these structures, the order of nodes should not influence the computation of predicted values, a property referred to as “permutation invariance.” Various approaches are used to implement GNNs, but one of the most prominent paradigms is based on “message passing.”
@@
where $\bigoplus$ represents a differentiable, permutation-invariant function, such as summation, averaging, or maximum operation, while $\gamma$ and $\phi$ denote differentiable functions, often implemented as neural network layers. In Equation 2, the model updates the node information using both existing information (from the previous layer) and information shared among neighboring nodes (messages), which is computed as in Equation 1. The latter equation pertains to the aggregation of messages from neighboring nodes.
@@
Our problem can be framed as the selection of models, specifically the graph representations. MDL [12, 13] is a paradigm of model selection that attempts to find the model that minimizes the number of bits representing information. This information may include the complexity of models and the performance score of the model on given data.
In our work, we use MDL to study a model whose complexity is minimized while maintaining the prediction’s well-performed accuracy. It is the two-stage code description approach [14]. Specifically, it encodes the data $D$ with the description length $\mathcal{L}_{\mathcal{H}}(D)$ by first encoding a hypothesis $H$ from the set of given hypotheses $\mathcal{H}$ followed by encoding $D$ with the help of $H$. The objective is to minimize the total description length as follows:
@@
III Related Work
III-A Feature Interaction Modeling
@@
From the machine learning perspective, many models were proposed to model feature interactions, especially variants of factorization models inspired by FM [15]. FM relies on the inner product between embedding the feature-value, motivated by the factorization of matrices. Several variants of FM have been proposed to fill the gaps. FFM [16] takes into account the embedding of feature field (i.e. columns in a table of data). AFM [17] considers the weight of the interaction by adding attention coefficients to its model.
@@
The attention mechanisms [19] are a method designed to model the importance of different components. In AFM, the attention mechanism was applied because the authors believe that not all interactions should be treated equally likely with respect to prediction values. Attention allows the model to weigh feature interactions differently based on their importance. After introducing the attention mechanism to interaction tasks, many works use attention to model the interaction of features based on the factorization machine framework [20, 21, 22, 23, 24].
One crucial challenge of interactions in traditional machine learning models is the explainability of which features have interactions, although some of them outperform in terms of prediction performance. So we need an explicit representation that can correspond to pairwise interaction, where a graph is a representation that can fill this gap. In addition, GNN is a tool that can be used to learn information from graphs.
III-B Graph Neural Networks for Feature Interactions
GNNs are becoming more interesting for use in feature interaction problems. Since feature interactions can be seen as relationships between features, most of the works rely on feature graphs for each individual instance to represent the interactions via the aggregation of representations from neighboring nodes.
Motivated by the click-through rate (CTR) prediction, which is believed to influence the rate of clicking on a product, such as age and gender, many works attempted to model this prediction using GNN applied to feature graphs. To the best of our knowledge, Fi-GNN [9] is the first work to use GNN to model the interaction of features in the feature graph for the prediction of CTR. It applied the field-aware embedding layer to compute the latent representation of nodes of features before feeding the feature graph with the computed representations into a stack of message passing layers.
@@ III-B Graph Neural Networks for Feature Interactions
All mentioned works chronically improve the methods of feature interaction learning. However, no work provides information about the effect of the construction of feature graphs on the learning capacity. Most works utilize attention mechanisms or edge weights to let models learn graph structures via these parameters.
IV Problem Formulation: Pairwise Interaction Feature Graph Problem
A crucial open research question in graph representation for data without explicit connections is that What is a suitable graph structure? While current research highlighted the potential of feature graphs, a feature graph that is too dense, for example, a complete graph, may lead to learning issues, as shown and discussed in previous work [9, 28, 29] and in Section VII-E. Thus, the community’s common goal is to find an optimal feature graph structure for any particular prediction task. Consequently, we can formulate the “feature-graph-finding” problem as follows:
@@ IV Problem Formulation: Pairwise Interaction Feature Graph Problem
where $\text{GNN}(\ \cdot\ ;G)$ is a given GNN model trained for the feature graph $G$, and $\ell$ is an arbitrary but fixed loss function for the prediction task. The above task is not traceable as there are exponentially many possible graph structures ($\mathcal{O}(2^{n^{2}})$.) Many works attempted to solve this problem using soft learnable interactions [9, 26] or hard interactions through reinforcement learning [10]. However, they did not provide insight into the characteristics of a suitable feature graph structure related to the interaction between features.
@@ V-A Correspondence Problem
Multiple graphs can infer the same expression, considering the perspective of connected components. For example, the expression $x_{0}x_{1}x_{2}+x_{3}x_{4}+x_{5}+x_{6}$ corresponds to the graphs $E_{1}=\{\{0,1\},\{0,2\},\{1,2\},\{3,4\}\}$ and also $E_{2}=\{\{0,1\},\{1,2\},\{3,4\}\}$. Consequently, we use equivalence classes of graphs (based on connectedness) instead of individual graphs. To this end, we define an equivalence class of feature graphs.
@@
V-B MDL for graph representation of pairwise interaction
@@
Problem 1 MDL Pairwise Interaction Graph Problem
@@
In this problem, we denote $\mathcal{L}(S,f_{G})=\mathcal{L}(f_{G})+\mathcal{L}(S|f_{G})$, where
@@
Finally, we get the result that in aspect of MDL, complete graph and null graph are worse than any graph containing all interaction edges stated as follows:
@@
The theoretical results shown in Section V-A are just to ensure us that we can confide in learning feature graph structures to infer an expression containing feature interactions. In other words, given any function containing feature interactions of at most one occurrence for each feature, there always exist feature graph structures that can infer backward to the function and vice versa. Therefore, to find such graph structures, we need some insight into graph construction related to feature interactions.
Consequently, we reformulate these observations into more general arguments. Generally speaking from empirical results in Section VII-B, it is better to maintain all the edges of the interaction in the feature graph to maximize the learning interaction in a given dataset. Also, it is better to prune non-interaction edges out as much as possible to avoid fitting noise information. Finally, these results are supported by Proposition 5.
In addition, we also show the empirical result about using complete feature graphs to learn from larger datasets in Subsection VII-E. Although the result in Subsection VII-B cannot infer the limitation of complete graph obviously, the result in Subsection VII-E reveals that the complete feature graphs are worse when datasets are larger in feature size. Finally, we can show that, in the manner of MDL which is a framework of model selection, the complete feature graph is worse than any feature graphs containing all interaction edges. This tells us that it is not a good idea to use the complete feature graph to learn from datasets. Not only learning hidden noise, but it also faces computation and memory issues. The following section will focus on the empirical results in more detail.
VI Methodology
@@
Our experiment provides empirical evidence regarding the relationship between feature graph structures and dataset pairwise interactions. We train and evaluate several feature graph structures for each dataset and report the results. However, due to the combinatorial number of possible feature graph structures and the stochastic nature of the learning algorithm, we adopt a strategy of generating random edge sets to explore various graph configurations, as illustrated in Figure 2.
VI-A Datasets
Real-world datasets rarely provide information about interactions between features, which disables our analysis. To this end, we use synthesized datasets to control interaction characteristics in datasets such as the number of interaction pairs of features, the strength of interactions, or the kinds of interactions. We assume that there is only one mode of interaction per feature graph. If real-world datasets had many interaction modes, we could apply the weighted ensemble feature graphs proposed by Li et al. [24]
@@ VI-C GNN Architectures
As Fi-GNN does, this work adopts the idea from Fi-GNN, but with a simpler architecture. We use TransformerConv [30], which is a message-passing layer based on the attention mechanism, whose feature interaction learning capability is shown by many works. For the computation, we start from the initial node embedding $H^{(0)}$ defined in Equation 9. Then, a feature graph with the node embedding matrix is fed to TransformerConv with BatchNorm:
@@
where $H^{(l)}$ is an intermediate node embedding matrix at the layer $l$. After applying all message passing layers, we calculate the average pooling of node embeddings from the last GNN layer, and we use this vector as a graph embedding, which is fed into the projection layer to get a prediction:
@@
where $H^{(L)}$ is a node embedding matrix from the last GNN layer and $H^{(L)}[i]$ is the row $i$ of the matrix which corresponds to the node embedding of the node $i$ and $b$ is a biased term. Equation 11 is the calculation of the mean pooling layer, and Equation 12 is the final projection to calculate the prediction value.
Equation 10 is a bit different from the original work [30], which uses LayerNorm after aggregation of messages, similar to the original transformer in NLP. In the experiment’s first phase, we found that the GNN model with LayerNorm cannot capture interaction behavior while BatchNorm can.
In training the GNN model, we adopt a usual supervised learning approach, where the loss function measures the difference between the predicted value, $\hat{y}$, and the target value, $y$. Specifically, we use the mean squared error (MSE) loss, which is defined as